Emil Wauseka is buying a used car with a total purchase price including interest of $14, 987. he plans to make $3,000 down payment and to finance the rest for 5 years. about how much will his monthly payments be?
Answer by JulietG(1812)

Assuming no points and no additional interest...

$14,987 - $3,000 = $11,987
$11,987 / 60 months (5 years = 60 months) = $199.78 per month

Check:
199.78*60 = 11,986.80 + 3000 = 14,986.80
(note: his last payment will be .20 higher)